OpenStudy (anonymous):

3x + 5y = 25 2x − 7y = −4

jhonyy9 (jhonyy9):

3x+5y=25 multiply by 2 2x-7y= -4 multiply by -3 6x+10y=50 -6x+21y=12 ------------ 0 31y=62 y=62/31 y=2 2x-7*2= -4 2x-14=-4 2x=10 x=10/2 x=5 x=5 y=2